# Neural Concept Verifier: Scaling Prover-Verifier Games via Concept Encodings

### v2

@@ 3.5 Training and Inference
  • Merlin’s loss:$L_{M}=CE(A(\mathcal{S}),y)$, where $\mathcal{S}$ is the sparse concept subset selected by Merlin. This loss encourages Arthur to classify correctly based on Merlin’s input.
  • Morgana’s loss:$L_{\widehat{M}}=CE(A(\widehat{\mathcal{S}}),y)$, where $\widehat{\mathcal{S}}$ is Morgana’s adversarial concept subset. Here, the loss is interpreted as the classifier’s inability to be misled by deceptive inputs [^1] In practice, the CE loss of Morgana is a slightly modified CE loss, cf. Wäldchen et al., (2024)..
@@
Overall, by integrating concept-extractor modules and leveraging the Merlin–Arthur framework, NCV emphasizes faithfulness and interpretability while preserving nonlinear modeling capabilities, and shifts the min–max optimization into a lower-dimensional concept space, improving efficiency, scalability, and stability in high-dimensional settings—a common challenge in min–max optimization for deep learning (Mescheder et al.,, 2018; Nagarajan and Kolter,, 2017).
@@ 4 Experimental Evaluations
In this section, we present a comprehensive evaluation of Neural Concept Verifier (NCV) on both synthetic and real-world high-dimensional image datasets. We evaluate based on two instantiations of NCV that utilize different concept extractors: a CLIP-based extractor and the Neural Concept Binder (NCB). We assess predictive performance and interpretability across multiple datasets, compare against established baselines, including SpLiCE and pixel-based Merlin-Arthur Classifiers, and examine scalability and robustness against shortcut learning.
Our evaluation is structured around the following research questions: (Q1) Does shifting Prover-Verifier Games (PVGs) to concept-encodings via NCV lead to performative classifiers on high-dimensional images (i.e., high completeness and soundness)? (Q2) Does NCV reduce the “interpretability-accuracy gap” in the context of CBMs? (Q3) Does NCV allow for more detailed explanations over pixel-based PVGs? Finally, (Q4) Can training via NCV reduce shortcut learning?
@@ 4.1 Experimental Setup
We investigate NCV in the context of the CLEVR-Hans3 and CLEVR-Hans7 (Stammer et al.,, 2021) datasets, which are synthetic benchmarks derived from CLEVR (Johnson et al.,, 2017) that capture complex object compositions and include visual shortcuts. CLEVR-Hans3 features three object classes, while CLEVR-Hans7 increases the task complexity to seven object classes. Both datasets consist of images rendered at a resolution of $128\!\times\!128$ pixels. The original datasets contain shortcuts from attributes to class labels in their training and validation sets, which are not present in the test set, making them well-suited for studying shortcut behavior. For example, in CLEVR-Hans3, a class may be defined by object shape and size, but during training, irrelevant features like color (e.g., gray cubes) are spuriously correlated with the label. Models relying on such shortcuts can fail under the decorrelated test distribution. We first evaluate our results on non-confounded versions of these datasets, where feature distributions are consistent across training, validation, and test splits. We return to the original confounded datasets when investigating shortcut mitigation.
Further, to assess scalability and generalization to natural image data, we evaluate on the widely-used benchmarks ImageNet-1k (Deng et al.,, 2009) and CIFAR-100 (Krizhevsky,, 2009). ImageNet-1k consists of over 1.2 million high-resolution images across 1,000 classes, typically resized to $224\!\times\!224$ pixels for training. CIFAR-100 provides 60,000 low-resolution images of $32\!\times\!32$ pixels across 100 fine-grained categories, representing a more compact but still challenging dataset to evaluate generalization.
4.1.2 Baselines
We compare our proposed framework against several representative baselines. Training parameters and implementation details for all baselines are provided in Suppl.  A.
Black-box Classifier: ResNet-18. We first consider a standard ResNet-18 model (He et al.,, 2016) trained end-to-end on raw image inputs (denoted as ResNet-18). This serves as a high-performance but non-interpretable baseline model.
Merlin-Arthur Classifier on Pixel-Space (Pixel-MAC). We further evaluate an instantiation of a Prover-Verifier Game, as a pixel-based Merlin-Arthur Classifier (Wäldchen et al.,, 2024) (denoted as Pixel-MAC). Hereby, the verifier is initialized from a pretrained ResNet-18 model, whereas both the cooperative prover (Merlin) and adversarial prover (Morgana) are implemented as U-Net models (Ronneberger et al.,, 2015) that output continuous-valued feature importance masks over the input image. These masks are subsequently discretized using a Top-$k$ selection to define the subset of features visible to the verifier. All agents are jointly fine-tuned within the prover-verifier setup. The explanations (i.e., certificates) in this setup represent masks on the original pixel space. For a detailed description of the general Pixel-MAC workflow, we refer to (Wäldchen et al.,, 2024).
Concept Bottleneck Model with Linear Classifier (CBM). Lastly, we compare NCV to a vanilla CBM setup in which the final classifier represents a linear model (denoted as CBM). In these baselines, a concept extractor first generates a vector of interpretable concept features from the input image. The resulting concept features are then passed to a linear classifier trained to predict the target class. We leverage NCB (Stammer et al., 2024b, ) as the concept extractor for evaluations on CLEVR-Hans and SpLiCE (Bhalla et al.,, 2024) as the concept extractor for evaluations on CIFAR-100 and ImageNet-1k.
4.1.3 NCV Model Instantiations
NCB-based Neural Concept Verifier. As detailed in Section  3.3, for evaluations on CLEVR-Hans3 and CLEVR-Hans7, we utilize NCB (Stammer et al., 2024b, ) to extract unsupervised, discrete concepts. Specifically, we use NCB models that were pretrained on the CLEVR (Johnson et al.,, 2017) dataset. NCB’s resulting concept encodings are structured as a set of object-centric slots. To handle this structure, we employ a permutation-invariant Set Transformer (Lee et al.,, 2019) as the verifier (Arthur). The provers (Merlin and Morgana) are implemented as independent neural networks, each parameterized by a Set Transformer. They receive the full concept slot matrix as input and output a sparse selection mask with 12 nonzero entries, indicating the active concept blocks provided to the verifier. All components are jointly trained using the Adam optimizer. Further details and ablation studies can be found in Suppl.  B.
CLIP-based Neural Concept Verifier. For evaluations on ImageNet-1k and CIFAR-100, we employ an instantiation of NCV with a CLIP-based concept extractor (Radford et al.,, 2021). We follow the approach of SpLiCE (Bhalla et al.,, 2024), which computes image-to-text similarity scores between CLIP image features and a predefined vocabulary of candidate concepts. However, where SpLiCE performs a per-sample optimization to obtain a sparse activation of the concepts, our NCV instantiation retains the full concept activation vector (denoted as CLIP-Sim in the following) and delegates the selection of relevant concepts to the subsequent prover agent in our NCV framework. This avoids expensive optimization at inference time and allows our approach to scale efficiently while preserving the ability to select meaningful concepts for interpretable classification. The verifier, Merlin and Morgana are all implemented as two-layer multilayer perceptron (MLP). The two provers receive the full concept activation vector as input and output a sparse selection mask of $32$ concepts per example. All components are respectively trained using the Adam optimizer. Further details and ablation studies can be found in Suppl.  C.
4.1.4 Metrics.
All methods are evaluated in terms of completeness and soundness (where applicable), as described in Section  3.4. Note that completeness is equivalent to standard classification accuracy in our setting, as it evaluates the model’s accuracy under cooperative feature selection, whereas soundness describes the robustness against adversarial prover selections. We evaluate with 20 random seeds for the CLEVR-Hans evaluations and 10 seeds for ImageNet-1k and CIFAR-100. We report the average performance and standard deviation over these seeds in all tables. Furthermore, when evaluating shortcut learning in the context of the CLEVR-Hans datasets, we report the validation–test gap, defined as the difference between validation accuracy on a confounded split and test accuracy on a non-confounded split. A smaller gap indicates improved generalization and reduced reliance on spurious correlations.
@@ 4.1.3 NCV Model Instantiations
Table 1: NCV delivers high predictive performance and robustness through verifiable, concept-based reasoning. We report completeness and soundness scores for ResNet-18, Pixel-MAC, CBM, and NCV across synthetic (CLEVR-Hans3, CLEVR-Hans7) and real-world (CIFAR-100, ImageNet-1k) datasets. NCV consistently matches or outperforms baselines in completeness while offering strong soundness guarantees.
Model Feature Space Completeness (Accuracy) Soundness (Robustness)
CLEVR-Hans3
[…]
CBM NCB 95.44± 0.0895.44plus-or-minus0.0895.$4$ 0.08 n/a
NCV (ours) NCB 98.92± $\boldsymbol{98.92\mbox{\scriptsize$\pm\,0.32$}}$ 100.00± $\boldsymbol{100.00\mbox{\scriptsize$\pm\,0.00$}}$
CLEVR-Hans7
[…]
NCV (ours) NCB 97.89± $\boldsymbol{97.89\mbox{\scriptsize$\pm\,0.31$}}$ 100.00± $\boldsymbol{100.00\mbox{\scriptsize$\pm\,0.00$}}$
CIFAR-100
ResNet-18 pixel space 79.73 ± $\pm\,0.36$ n/a
[…]
CBM SpLiCE 75.42 ± $\pm\,0.04$ n/a
NCV (ours) CLIP-Sim 83.32± $\boldsymbol{83.32\mbox{\scriptsize$\pm\,0.28$}}$ 99.99± $\boldsymbol{99.99\mbox{\scriptsize$\pm\,0.01$}}$
ImageNet
ResNet-18 pixel space 66.16± $\pm\,0.41$ n/a
[…]
CBM SpLiCE 68.59± $\boldsymbol{68.59\mbox{\scriptsize$\pm\,0.01$}}$ n/a
[…]
@@ 4.2 Evaluations
Scalability to High Dimensions (Q1).
In our first evaluation, we examine whether shifting the Prover–Verifier Game (PVG) to concept encodings enables NCV to scale to high-dimensional image domains while achieving strong performance in terms of completeness and soundness. We hereby compare NCV against two key baselines: (1) a black-box ResNet-18 classifier, and (2) Pixel-MAC, a nonlinear PVG model operating in raw pixel space. Table  1 summarizes results across synthetic (CLEVR-Hans3, CLEVR-Hans7) and real-world (CIFAR-100, ImageNet-1k) benchmarks. Each model’s feature space is indicated for clarity. On the synthetic CLEVR-Hans benchmarks, we observe that NCV consistently achieves the highest completeness scores, surpassing Pixel-MAC and even ResNet-18 on CLEVR-Hans3, while also attaining perfect soundness. This demonstrates that NCV not only matches or exceeds the performance of strong black-box classifiers but also maintains faithful and certifiable decision-making. Pixel-MAC performs well in these settings but falls slightly short in completeness and cannot match NCV’s zero-error soundness.
On the more challenging real-world datasets CIFAR-100 and ImageNet-1k, Pixel-MAC either fails entirely or performs poorly. In contrast, NCV successfully scales to these datasets, achieving superior completeness and near-perfect soundness without incurring test-time optimization overhead. Notably, NCV surpasses ResNet-18 even in raw accuracy, providing both higher predictive performance while retaining interpretability. In summary, NCV generalizes well across domains: it scales beyond the limitations of pixel-based PVGs, delivers competitive accuracy even on large-scale and complex datasets, and retains soundness throughout. These findings affirm that shifting PVGs to concept space enables interpretable and verifiable classifiers that remain performant and scalable in high-dimensional visual environments. We therefore answer Q1 affirmatively.
@@
In Table  1, we further examine whether NCV can overcome a central limitation of standard Concept Bottleneck Models (CBMs): the interpretability–accuracy gap resulting from their use of constrained linear classifiers. Across all evaluated datasets, NCV consistently narrows—and in some cases closes—this gap, while maintaining formal verifiability. On CLEVR-Hans3, CBM trails the opaque ResNet-18 by over 2 percentage points in completeness, whereas NCV matches or exceeds the ResNet’s performance while retaining perfect soundness. The benefit is even more pronounced on CLEVR-Hans7: CBM underperforms ResNet-18 by nearly 10 percentage points, while NCV narrows the gap to just 1 percentage point and again delivers flawless soundness. This trend persists on real-world datasets. On CIFAR-100, NCV outperforms CBM by nearly 8 percentage points in completeness and even slightly exceeds ResNet-18’s performance, while providing robust and interpretable predictions. On ImageNet-1k, CBM slightly outperforms NCV in raw completeness by a small margin. Overall, NCV consistently improves over linear CBMs in both accuracy and robustness, especially on tasks requiring complex concept reasoning. At the same time, it matches or even surpasses the accuracy of the ResNet-18 model, demonstrating that interpretable, concept-level reasoning via Prover–Verifier Games can deliver competitive performance without sacrificing transparency or verifiability. We therefore answer Q2 affirmatively.
Higher Detailed Explanations (Q3)
@@
We next investigate the resulting explanations produced by our NCV framework, with a focus on interpretability. Since our goal is to improve the explanatory clarity of Prover–Verifier Games, we compare against pixel-level MAC explanations, which serve as a natural baseline for interpretability within the PVG framework. Figure  3 illustrates a qualitative example from the CLEVR-Hans3 dataset (cf. Figure  4 for more examples), both from the Pixel-MAC setup and NCV. Notably, under Pixel-MAC (see top panel of Figure  3) the Prover–Verifier setup operates directly on pixels, yielding broad, diffuse masks that often cover entire objects or irrelevant background regions, arguably providing limited insight regarding which exact features drive the verifier’s final decision. These patterns also reflect typical failure modes of Pixel-MAC despite high completeness and soundness. In contrast, NCV leverages its internal concept encodings $\mathbf{c}$ to pinpoint the precise attributes relevant for a class decision (see bottom panel of Figure  3), e.g., a “small metal cube” and “small metal sphere”. [^2] It is important to note that this object-level specification is dependent on the utilized concept extractor. NCB, by design allows to receive object-specific concept explanations. However, this is by default not possible via the CLIP-based concept extractor. In any case, NCV allows for receiving explanations on the level of its concept-encodings. Overall, these examples highlight that NCV offers substantially more detailed and understandable explanations than pixel-based PVG baselines, highlighting that concept-level PVGs yield verifiable decisions that are potentially aligned with underlying logical rules. This leads us to answer Q3 affirmatively.
Figure 3: Comparison of explanations from NCV vs. Pixel-MAC for a CLEVR-Hans3 image (class label two) with GT class rule: small metal cube and small metal sphere. (a) Merlin–Arthur training on pixel space yields uninformative masks. (b) MAC on concept encodings via NCV provides clear explanations by highlighting object features corresponding to the class rule. The single-object images are reconstructions from the respective slots selected by Merlin (prover).

### v3

@@ 3.5 Training and Inference
[…]
@@
Overall, by integrating concept-extractor modules and leveraging the Merlin–Arthur framework, NCV emphasizes faithfulness [^2] Our notion of ‘faithfulness’ is defined in Sec.  A.3. and interpretability while preserving nonlinear modeling capabilities, and shifts the min–max optimization into a lower-dimensional concept space, improving efficiency, scalability, and stability in high-dimensional settings, a common challenge in min–max optimization for deep learning (Mescheder et al., 2018; Nagarajan and Kolter, 2017).
@@ 4 Experimental Evaluations
In this section, we present a comprehensive evaluation of Neural Concept Verifier (NCV) on both synthetic and real-world high-dimensional image datasets. We evaluate based on two instantiations of NCV that utilize different concept extractors: a CLIP-based extractor and the Neural Concept Binder (NCB) (Stammer et al., 2024b). We assess predictive performance and interpretability across multiple datasets, compare against several baselines and examine scalability as well as robustness against shortcut learning.
Our evaluation is structured around the following research questions: (Q1) Does shifting Prover-Verifier Games (PVGs) to concept-encodings via NCV lead to performative classifiers (i.e., high completeness and soundness) on high-dimensional synthetic and real-world images? (Q2) Does NCV reduce the “interpretability-accuracy gap” in the context of CBMs? (Q3) Does NCV allow for more detailed explanations over pixel-based PVGs? Finally, (Q4) Can training via NCV reduce shortcut learning?
@@ 4.1 Experimental Setup
Datasets. We investigate NCV on CLEVR-Hans3 and CLEVR-Hans7 (Stammer et al., 2021), synthetic benchmarks derived from CLEVR (Johnson et al., 2017) that capture complex object compositions and include visual shortcuts. CLEVR-Hans3 features three compositional classes, while CLEVR-Hans7 increases the complexity to seven, with all images rendered at $128\!\times\!128$ pixels. The training and validation sets contain spurious correlations between attributes and labels (e.g., gray cubes linked to a specific class), which are absent in the test set, making the datasets well-suited for studying shortcut behavior. Models that exploit such correlations often fail under the decorrelated test distribution. We first report results on non-confounded versions of these datasets, where feature distributions are consistent across splits, and later return to the confounded versions for shortcut mitigation. To assess scalability and generalization to natural images, we additionally evaluate on ImageNet-1k (Deng et al., 2009) with 1.2M high-resolution images across 1,000 classes (resized to $224\!\times\!224$ pixels), and on CIFAR-100 (Krizhevsky, 2009) with 60,000 low-resolution $32\!\times\!32$ images across 100 fine-grained categories. Lastly, we perform experiments on COCOLogic (Steinmann et al., 2025), a recent benchmark combining real-world images with complex, compositional class rules.
Baseline Models. We compare our framework against several representative baselines, with training details provided in Suppl.  C. As a strong but non-interpretable baseline, we use a standard ResNet-18 (He et al., 2016) for evaluations on CLEVR-Hans, and a ResNet-50 for CIFAR-100, COCOLogic and ImageNet-1k, each trained end-to-end on raw images. We further evaluate a pixel-based MAC (Wäldchen et al., 2024) (denoted as Pixel-MAC), an instantiation of the Prover-Verifier Game in which the verifier is initialized from a pretrained ResNet-18, while both provers (Merlin and Morgana) are U-Net models (Ronneberger et al., 2015) that output continuous feature-importance masks over the input image. These masks are discretized using Top-$k$ selection to define the features visible to the verifier, and all agents are jointly fine-tuned; the resulting explanations (i.e., certificates) correspond to masks in pixel space (see (Wäldchen et al., 2024) for further details). We also compare to a vanilla Concept Bottleneck Model (Koh et al., 2020) (denoted as CBM), where a linear classifier predicts from concept features extracted by either NCB (Stammer et al., 2024b) for CLEVR-Hans or SpLiCE (Bhalla et al., 2024) for CIFAR-100, ImageNet-1k and COCOLogic. In addition, we include a nonlinear CBM variant that replaces the linear classifier head with a nonlinear, non-interpretable head while operating on the same concept encodings, isolating the effect of a more expressive concept-level predictor.
@@
Table 1: NCV delivers high predictive performance and soundness through verifiable, concept-based reasoning evaluated via completeness and soundness. We report completeness and soundness scores for ResNet, Pixel-MAC, nonlinear and linear CBMs, and NCV across synthetic (CLEVR-Hans3, CLEVR-Hans7) and real-world (CIFAR-100, ImageNet-1k, COCOLogic) datasets. NCV matches or outperforms baselines in completeness in most settings, while offering strong soundness guarantees.
Model Feature Space Completeness (Accuracy) Soundness (Robustness) Completeness (Accuracy) Soundness (Robustness) Completeness (Accuracy) Soundness (Robustness)
CIFAR-100 ImageNet-1k COCOLogic
ResNet-50 pixel space $81.45\mbox{\scriptsize$\pm\,0.60$}$ n/a $76.01\mbox{\scriptsize$\pm\,0.02$}$ n/a $65.80\mbox{\scriptsize$\pm\,3.41$}$ n/a
CBM (nonlin.) SpLiCE $79.29\mbox{\scriptsize$\pm\,0.42$}$ n/a $69.02\mbox{\scriptsize$\pm\,0.38$}$ n/a $70.09\mbox{\scriptsize$\pm\,0.56$}$ n/a
Pixel-MAC pixel space $15.27\mbox{\scriptsize$\pm\,4.78$}$ $96.31\mbox{\scriptsize$\pm\,4.12$}$ $35.06\mbox{\scriptsize$\pm\,3.20$}$ $99.65\mbox{\scriptsize$\pm\,0.26$}$ $42.57\mbox{\scriptsize$\pm\,3.13$}$ $97.70\mbox{\scriptsize$\pm\,0.61$}$
CBM SpLiCE $75.42\mbox{\scriptsize$\pm\,0.04$}$ n/a $\bm{68.59\mbox{\scriptsize$\pm\,0.01$}}$ n/a $58.84\mbox{\scriptsize$\pm\,0.09$}$ n/a
NCV (ours) CLIP-Sim $\bm{83.32\mbox{\scriptsize$\pm\,0.28$}}$ $\bm{99.99\mbox{\scriptsize$\pm\,0.01$}}$ $67.04\mbox{\scriptsize$\pm\,0.16$}$ $\bm{99.94\mbox{\scriptsize$\pm\,0.02$}}$ $\bm{75.42\mbox{\scriptsize$\pm\,3.21$}}$ $\bm{97.87\mbox{\scriptsize$\pm\,0.47$}}$
CLEVR-Hans3 CLEVR-Hans7
ResNet-18 pixel space $97.87\mbox{\scriptsize$\pm\,0.24$}$ n/a $98.71\mbox{\scriptsize$\pm\,0.24$}$ n/a
CBM (nonlin.) NCB $98.13\mbox{\scriptsize$\pm\,0.37$}$ n/a $97.83\mbox{\scriptsize$\pm\,0.25$}$ n/a
Pixel-MAC pixel space $6.59\mbox{\scriptsize$\pm\,0.72$} 99$.99± $99\mbox{\scriptsize$\pm\,0.01$}$ $7.61\mbox{\scriptsize$\pm\,0.38$} 99$$9.88\mbox{\scriptsize$\pm\,0.28$}$
CBM NCB $95.44\mbox{\scriptsize$\pm\,0.08$}$ n/a $89.12\mbox{\scriptsize$\pm\,0.12$}$ n/a
NCV (ours) NCB $\bm{98.92\mbox{\scriptsize$\pm\,0.32$}}$ $\bm{100.00\mbox{\scriptsize$\pm\,0.00$}}$ $\bm{97.89\mbox{\scriptsize$\pm\,0.31$}}$ $\bm{100.00\mbox{\scriptsize$\pm\,0.00$}}$
NCV Instantiations. For CLEVR-Hans3 and CLEVR-Hans7, we instantiate NCV with NCB (Stammer et al., 2024b) as the concept extractor, using models pretrained on CLEVR (Johnson et al., 2017). A permutation-invariant Set Transformer (Lee et al., 2019) serves as the verifier (Arthur) to process the unordered NCB encodings. The provers (Merlin and Morgana) are independent Set Transformers that take the full concept-slot encodings as input and output a sparse mask of 12 active concepts for the verifier. All agents are jointly trained with the Adam optimizer (Kingma and Ba, 2014). Further details and ablations are provided in Suppl.  D. For ImageNet-1k, CIFAR-100 and COCOLogic, we use a CLIP-based concept extractor (Radford et al., 2021), following the approach of SpLiCE (Bhalla et al., 2024) to compute image–text similarity scores. All concepts are drawn from a fixed textual vocabulary derived from LAION captions (Schuhmann et al., 2021) and embedded using CLIP (see Suppl.  E). Unlike SpLiCE, which performs per-sample optimization, our method (denoted as CLIP-Sim) retains the full activation vector and delegates concept selection to the provers, avoiding expensive inference-time optimization and enabling scalability. Here, the verifier and both provers are two-layer MLPs; the provers output sparse masks of 32 concepts per example. All agents are trained with the Adam optimizer. Additional details and ablations are provided in Suppl.  D and Suppl.  E including the effect of varying the mask size and the weighting parameter $\gamma$.
Metrics. All methods are evaluated for completeness and, where applicable, soundness (cf. Sec.  3.4). We use 20 random seeds for CLEVR-Hans and 10 for ImageNet-1k, CIFAR-100 and COCOLogic, reporting mean and standard deviation across all seeds. For shortcut learning evaluations with CLEVR-Hans, we additionally report a separate shortcut robustness metric: the validation-test gap, i.e., the difference between validation accuracy on data with shortcuts and test accuracy on a split without these shortcuts.
@@ 4.2 Evaluations
Scaling PVGs to High Dimensions (Q1).
In our first evaluation, we examine whether shifting the Prover–Verifier Game (PVG) to concept encodings enables NCV to scale to high-dimensional image domains while achieving strong performance in terms of completeness and soundness. We hereby compare NCV against two key baselines: (1) a black-box ResNet classifier (ResNet-18 for CLEVR-Hans and ResNet-50 for CIFAR-100, ImageNet-1k and COCOLogic), and (2) Pixel-MAC, a nonlinear PVG model operating in raw pixel space. Tab.  1 summarizes results across synthetic (CLEVR-Hans3, CLEVR-Hans7) and real-world (CIFAR-100, ImageNet-1k, COCOLogic) benchmarks. Each model’s feature space is indicated for clarity. On the synthetic CLEVR-Hans benchmarks, we observe that NCV consistently achieves the highest completeness scores, surpassing Pixel-MAC and even ResNet-18 on CLEVR-Hans3, while also attaining perfect soundness. This demonstrates that NCV not only matches or exceeds the performance of strong black-box classifiers but also certifiable decision-making. Pixel-MAC performs well in these settings but falls slightly short in completeness and cannot match NCV’s zero-error soundness.
On the more challenging real-world datasets, Pixel-MAC either fails entirely or performs poorly. In contrast, NCV successfully scales to these datasets, achieving superior completeness and near-perfect soundness. Notably, NCV surpasses ResNet-50 even in raw accuracy for CIFAR-100 and COCOLogic, providing both higher predictive performance while retaining interpretability. In summary, NCV generalizes well across domains: it scales beyond the limitations of pixel-based PVGs, delivers competitive accuracy even on large-scale and complex datasets, and retains soundness throughout. Overall, these findings affirm that shifting PVGs to concept space enables interpretable classifiers whose decisions can be evaluated via completeness and soundness, while remaining performant and scalable in high-dimensional synthetic and real-world visual environments. We therefore answer Q1 affirmatively.
Figure 3: Comparison of explanations from NCV vs. Pixel-MAC. (top) Merlin–Arthur training on pixel space yields uninformative masks. (bottom) MAC on a concept-space via NCV translates into combinations of high-level concepts and, in turn, in an interpretable prediction.
@@
In Tab.  1, we further examine whether NCV can overcome a central limitation of a majority of Concept Bottleneck Models (CBMs): the interpretability–accuracy gap resulting from their use of constrained linear classifiers (cf. Suppl.  B for a discussion). We observe that NCV consistently narrows or closes this gap while maintaining high completeness and soundness. Specifically, on CLEVR-Hans3, NCV ($\sim$99%) exceeds the linear CBM ($\sim$95%) and both the opaque ResNet-18 and nonlinear CBM baselines ($\sim$98%) with perfect soundness. On CLEVR-Hans7, NCV narrows the 10-point gap left by linear CBMs to match the performance of its nonlinear counterparts. This trend persists on real-world datasets. On CIFAR-100, NCV outperforms both the linear and nonlinear CBMs and even slightly exceeds ResNet-50’s performance. This is even more pronounced on COCOLogic, where NCV outperforms all other models by a large margin. As the additional benefits of a nonlinear classifier are quite small on ImageNet-1k, the additional training complexity of NCV results in a slightly worse performance compared to the CBMs there. Overall, NCV improves over linear CBMs in accuracy, especially on tasks requiring complex concept reasoning such as COCOLogic. At the same time, NCV can match or even surpass the accuracy of opaque ResNet models, demonstrating that interpretable, concept-level classification via Prover–Verifier Games can deliver competitive performance. Additional comparisons with Deep Concept Reasoning (DCR; Barbiero et al. 2023) using the same concept vocabulary are reported in the supplementary material, highlighting that NCV can also outperform CBMs with nonlinear, yet interpretable classifiers on large-scale datasets (cf. Tab.  4 in Suppl.  C). We therefore answer Q2 affirmatively.
More Detailed Explanations (Q3)
@@
We next investigate the resulting explanations produced by NCV, with a focus on explanatory clarity. Since our goal is to improve over classic vision-based Prover–Verifier Games, we compare against pixel-level MAC explanations. Fig.  3 illustrates a qualitative example from both the CLEVR‐Hans3 (cf. Fig.  4 for more examples) and ImageNet-1k datasets. Notably, under Pixel‐MAC, the Prover–Verifier setup operates directly on pixels, yielding broad, diffuse explanation masks that often cover entire objects or irrelevant background regions, arguably providing limited insight regarding which exact features drive the verifier’s final decision. In contrast, NCV leverages its internal concept encodings to isolate sparse, high-level concepts that are consistently associated with a class decision under the prover–verifier interaction, recovering the class rule for CLEVR-Hans (i.e., small cube and small metal sphere) and providing a meaningful concept explanation for the coffee-mug class of ImageNet-1k. [^3] The availability of object-level concepts in NCV depends on the underlying concept extractor. For CLEVR-Hans, we use NCB, which provides such object-based explanations. For clarity, we only show the five most frequent concepts (from a total mask size of 32) across 32 samples for ImageNet-1k. All concepts are part of the CLIP/SpLiCE-based concept vocabulary that is derived from LAION captions (Schuhmann et al., 2021) (see Suppl.  E for more details).
Overall, these examples highlight that NCV offers higher-level, semantically meaningful explanations rather than fine-grained pixel masks, and that concept-level PVGs yield interpretable decisions whose supporting concept subsets can be evaluated even for complex, high-dimensional data. This leads us to answer Q3 affirmatively.
Table 2: Shortcut robustness on CLEVR-Hans3 and CLEVR-Hans7. We report validation accuracy on a shortcut-confounded split, test accuracy on a clean split, and the resulting validation–test gap (“shortcut robustness”, lower is better) across models trained with varying amounts of clean data. When any amount of clean data is available, NCV shows improvements compared to training without PVGs.
Ratio Non-Conf. (Samples) Model CLEVR-Hans3 CLEVR-Hans7
Val Acc (w/ shortcut) Test Acc (w/o shortcut) Val-Test Gap (↓) Val Acc (w/ shortcut) Test Acc (w/o shortcut) Val-Test Gap (↓)
0% CBM (lin.) $95.65\mbox{\scriptsize$\pm\,0.09$}$ $90.54\mbox{\scriptsize$\pm\,0.09$}$ $5.11$ $90.37\mbox{\scriptsize$\pm\,0.10$}$ $85.27\mbox{\scriptsize$\pm\,0.15$}$ $\mathbf{5.10}$
CBM (non-lin.) $98.70\mbox{\scriptsize$\pm\,0.32$}$ $\mathbf{95.04}\mbox{\scriptsize$\pm\,0.96$}$ $\mathbf{3.66}$ $98.09\mbox{\scriptsize$\pm\,0.24$}$ $90.69\mbox{\scriptsize$\pm\,1.17$}$ $7.40$
NCV $\mathbf{99.44}\mbox{\scriptsize$\pm\,0.15$}$ $94.21\mbox{\scriptsize$\pm\,1.41$}$ $5.23$ $\mathbf{98.38}\mbox{\scriptsize$\pm\,0.18$}$ $\mathbf{92.23}\mbox{\scriptsize$\pm\,0.67$}$ $6.15$
1% (105) CBM (lin.) $96.28\mbox{\scriptsize$\pm\,0.16$}$ $91.03\mbox{\scriptsize$\pm\,0.31$}$ $5.25$ $90.74\mbox{\scriptsize$\pm\,0.12$}$ $85.41\mbox{\scriptsize$\pm\,0.17$}$ $5.33$
CBM (non-lin.) $99.10\mbox{\scriptsize$\pm\,0.27$}$ $94.84\mbox{\scriptsize$\pm\,0.98$}$ $4.26$ $98.17\mbox{\scriptsize$\pm\,0.17$}$ $92.65\mbox{\scriptsize$\pm\,1.31$}$ $5.52$
NCV $\mathbf{99.37}\mbox{\scriptsize$\pm\,0.18$}$ $\mathbf{97.11}\mbox{\scriptsize$\pm\,0.98$}$ $\mathbf{2.26}$ $\mathbf{98.19}\mbox{\scriptsize$\pm\,0.24$}$ $\mathbf{94.68}\mbox{\scriptsize$\pm\,0.64$}$ $\mathbf{3.51}$
5% (525) CBM (lin.) $95.38\mbox{\scriptsize$\pm\,0.37$}$ $93.34\mbox{\scriptsize$\pm\,0.51$}$ $2.04$ $90.37\mbox{\scriptsize$\pm\,0.15$}$ $86.37\mbox{\scriptsize$\pm\,0.18$}$ $4.00$
CBM (non-lin.) $98.41\mbox{\scriptsize$\pm\,0.55$}$ $96.13\mbox{\scriptsize$\pm\,0.71$}$ $2.28$ $98.32\mbox{\scriptsize$\pm\,0.22$}$ $95.19\mbox{\scriptsize$\pm\,0.80$}$ $3.13$
NCV $\mathbf{99.59}\mbox{\scriptsize$\pm\,0.19$}$ $\mathbf{98.88}\mbox{\scriptsize$\pm\,0.37$}$ $\mathbf{0.71}$ $\mathbf{98.47}\mbox{\scriptsize$\pm\,0.24$}$ $\mathbf{96.24}\mbox{\scriptsize$\pm\,0.71$}$ $\mathbf{2.23}$
20% (2100) CBM (lin.) $95.67\mbox{\scriptsize$\pm\,0.28$}$ $93.46\mbox{\scriptsize$\pm\,0.23$}$ $2.21$ $89.93\mbox{\scriptsize$\pm\,0.29$}$ $87.21\mbox{\scriptsize$\pm\,0.31$}$ $2.72$
CBM (non-lin.) $99.15\mbox{\scriptsize$\pm\,0.21$}$ $98.09\mbox{\scriptsize$\pm\,0.51$}$ $1.06$ $98.21\mbox{\scriptsize$\pm\,0.29$}$ $97.00\mbox{\scriptsize$\pm\,0.49$}$ $1.21$
NCV $\mathbf{99.37}\mbox{\scriptsize$\pm\,0.28$}$ $\mathbf{98.82}\mbox{\scriptsize$\pm\,0.67$}$ $\mathbf{0.55}$ $\mathbf{98.63}\mbox{\scriptsize$\pm\,0.13$}$ $\mathbf{97.74}\mbox{\scriptsize$\pm\,0.28$}$ $\mathbf{0.89}$